OpenStudy (anonymous):

which ordered pair is a solution of y=5x a.-2,10 b.-5,25 c.-3,15 d.-2,-10

OpenStudy (anonymous):

First You need to try substituting the options

OpenStudy (anonymous):

The correct answer will be equal on both sides of the equation

OpenStudy (texaschic101):

this is just a matter of substitution to see which one is correct I will do the first one.. y = 5x (-2,10)...x = -2 and y = 10 now we sub 10 = 5(-2) 10 = -10 ...this is incorrect and is not a solution. now sub in your points and find the one that is correct

OpenStudy (anonymous):

is it b?

OpenStudy (texaschic101):

when you sub in your points....does it make the equation true ?

OpenStudy (anonymous):

im a bit confused

OpenStudy (anonymous):

on substituting

OpenStudy (texaschic101):

your points are set up like this....(x,y) y = 5x (-5,25)....x = -5 and y = 25 now sub 25 = 5(-5) 25 = -25.....is this correct ?

OpenStudy (anonymous):

For substituting for this problem you are going to simply take the first number in the problem and replace it with the X and replace the second number for the y and then multiply the x number by five since they are coordinates.

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

so b is not the solution. Now try your other answer choices and find the one that is true

OpenStudy (anonymous):

oh it's D

OpenStudy (texaschic101):

BINGO...yes...you got it

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

wait

OpenStudy (texaschic101):

ok

OpenStudy (anonymous):

can you help me again...

OpenStudy (texaschic101):

I can try

OpenStudy (anonymous):

which ordered pair is a solution of -7x+2 a. 1,2 b. 8,-54 c.5,-35 d.1,-7

OpenStudy (anonymous):

now i was worming out this one and thinking it's c

OpenStudy (anonymous):

working*

OpenStudy (texaschic101):

do it the same way......oh...wait....this is not the entire equation...whats it equal to ?

OpenStudy (anonymous):

yeah it's a question mark next to the 2

OpenStudy (texaschic101):

I do not understand....it is not an equation if there is no equal sign. this is what would happen if I just subbed in the points... -7x + 2 (1,2)...x = 1 and y = 2 now sub -7(1) + 2 = -7 + 2 = -5 there is not even a y in your equation....I cannot solve this

OpenStudy (anonymous):

sorry it's y= -7x+2

OpenStudy (anonymous):

@texaschic101